Determination of a Chemical Formula
Write the following series in sigma notation and find the exact sun N7. 2 4 8 + 9 1 - 27
Answer #1

finn. S: 1-29 + 4/8 paha $= 1 +(-43) + (-42) + (-20).... -O X N=0 ©x (-43) S2X (-43) = -23 + (243)?+ (-20) 4.... - © - © S: -